Assume risk free rate of 5pr cent (with 360 days) per year calculate the spot price, So upon initiation of the contract.

Future price = 14,10

5 days remaining until the delivery

ANSWER:

F0,5 =14,10 = So * e(-0,05) (5/360) \Leftrightarrow S0= 14,09

How do you get 14,09 ?

Is e(-0,05) (5/360) = 0,9993058
Answer #1

Hi

In this formula a slight error is there,

F = S0*e^(rt)

S = F*e^(-rt)

S = 14.10*e^(-0.05*5/360)

S = 14.10* 0.9993

S = 14.09

Thanks